OpenStudy (anonymous):

One cubit foot holds 7.48 gallons of water, and one gallon of water weights 8.33 lbs. how much does a cubit foot of water weigh in lbs? in Ounces?

OpenStudy (lgbasallote):

\[1 \; \text{ft}^3 = 7.48 \; \text{gal}\] \[1 \; \text{gal} = 8.33 \; \text{lbs}\] you can use dimensional analysis \[\large \implies 1 \; \text{ft}^3 \times \frac{7.48 \; \text{gal}}{1 \; \text{ft}^3} \times \frac{8.33 \; \text{lbs}}{1 \; \text{gal}}\] \[\large \implies 1 \; \cancel{\text{ft}^3} \times \frac{7.48 \; \cancel{\text{gal}}}{1 \; \cancel{\text{ft}^3}} \times \frac{8.33 \; \text{lbs}}{1 \; \cancel{\text{gal}}}\] so just calculae 7.48 x 8.33 and you'll get your answer does that help?

OpenStudy (anonymous):

Yes- thank you! I just cannot get a grasp on this dimensional analysis. What is an easy way to learn this? Do you know of any free resources to learn more about this?

OpenStudy (lgbasallote):

hmm....i don't know any sites

OpenStudy (lgbasallote):

but what you need to know is you're aiming for cancellation

OpenStudy (anonymous):

It's all about multiplying the fractions so that the units you don't want cancel out and the unit you do want is left in the numerator.

OpenStudy (anonymous):

Thank you all!